Word problem - hard one - thanks for help
Okay
Making 3 different kinds of missiles for defense - each one takes a different combination of alloys - aluminum, plexiglas, and titanium
Missile 1 - needs 2 tons of alum, 2 tons of plex and 3 tons of titan
Missile 2 - 3 tons of alum, and 1 ton of titan
Missile 3 - 3 tons of plex and 1 ton alum
Need to produce 800 of type 1 650 of type 2 and 350 of type 3
How many tons of each alloy will we need???? HELP!?!?!?!?!?
Steps needed:
1. assign variables
2. write algebraic model
3. solve model
4. answer
thank you so much - lucyloo

Answer by stanbon(75887)

Making 3 different kinds of missiles for defense - each one takes a different combination of alloys - aluminum, plexiglas, and titanium
Missile 1 - needs 2 tons of alum, 2 tons of plex and 3 tons of titan
Missile 2 - 3 tons of alum, and 1 ton of titan
Missile 3 - 3 tons of plex and 1 ton alum
Need to produce 800 of type 1 650 of type 2 and 350 of type 3
How many tons of each alloy will we need?
---------------
Let "a" be amt of alum; "p" be amt of plex; "t" amt of titan
--------
Total alum = 800*2 + 650*3 + 350*1 = 3900 tons
Total plex = 800*2 + 350*3 = 2650 tons
Total titan = 800*3 + 650*1 = 3050 tons
=============
Comment: I doubt this is the problem you were asked to work.
Are you sure the 800, 650, and 350 are not respectively the
amts of alem, plex, and titan, rather than the number of
each missile type?
Cheers,
Stan H.